OpenStudy (1dfanatic):

Solve using an equation or a proportion: 85% of k is 68. What is k?

OpenStudy (anonymous):

85/100=68/k then cross multiply then solve.

OpenStudy (owlcoffee):

You can view "85%" as: \[\frac{ 85 }{ 100 }k\] or: \[0.85k\] I will work with the first one, we will traduce the problem into a mathematical equation wich can help us to solve it using the axioms and algebraic propeties that you were taught in your classes. \[\frac{ 85 }{ 100 }k=68\] So, solving: \[85k=6800\] \[k=\frac{ 6800 }{ 85 }\] if you do 6800/85, you'll get the result.

OpenStudy (anonymous):

so k=80

OpenStudy (1dfanatic):

Thank you @Crichards9100